OpenStudy (anonymous):

solve the system of equation what is the solution y=-x-12 8x-7y=9

OpenStudy (anonymous):

\[y=-x-12\]\[8x-7y=9\] \[8x-7(-x-12)=9\]\[8x+7x+84=9\]\[15x=-15\]\[x=-5\] So\[y=-(-5)-12\]\[y=-7\]